OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle.

OpenStudy (anonymous):

\[\tan \theta = \frac{ opposite }{ adjacent }\]Since the side that measures 41 feet is opposite of the 30 degree angle, it will be adjacent to the 60 degree angle. So you know tan(60), and you know that the adjacent side is 41. Now just solve \[\tan30=\frac{ x }{ 41 }\]

OpenStudy (anonymous):

how would I check my answer?

OpenStudy (anonymous):

Well, what answer did you get?

OpenStudy (anonymous):

Opps, when I wrote tan 30 up there I meant tan 60. I apologize for the confusion.

OpenStudy (anonymous):

i'm almost positive it isn't right..... ok.. hang on

OpenStudy (anonymous):

60=x/41 is my formula, right?

OpenStudy (anonymous):

tan 60 = x/41 It's a trigonometric function. tan(60) = Sqrt(3)

OpenStudy (anonymous):

1.37?

OpenStudy (anonymous):

No.\[\tan60=\frac{ x }{ 41 }\]\[41 \times \tan 60=x\]\[41 \times \sqrt{3}=x\]\[x \approx 71.0141\]

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes, I think so...

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

Find the length of the hypotenuse of the triangular lot.?

OpenStudy (anonymous):

Ok. I think these equations will help you out a lot:\[\sin A = \frac{ opposite }{ hypotenuse }\]\[\cos A = \frac{ adjacent }{ hypotenuse }\]\[\tan A = \frac{ opposite }{ adjacent }\] Since you want to know the hypotenuse, you'll want to use either sin or cos. It doesn't matter which you use, as long as you know which angle and side to use it on. But in your problem statement, you're given that the side that is 41 is opposite the 30 degree angle, so why not use sin(30)=41/x?

OpenStudy (anonymous):

82?

OpenStudy (anonymous):

Yep! :D

OpenStudy (anonymous):

thanks... one more....

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

sin30 =1/2 cos30=Sqrt(2)/2 tan30=Sqrt(3)/3 Can you put those into decimal forms?

OpenStudy (anonymous):

sin30=.5000 cos30=.8660 tan30=.5773

OpenStudy (anonymous):

Yes, almost. tan30=.5773503... So when rounding you would say tan30=.5774 Other than that, everything you have is correct. Sorry about the long delay, my internet went out for a few minutes.